OpenStudy (anonymous):

The length of a rectangle is 5/6 ft. The width is 3/8 ft. How much greater is the length of the rectangle than the width? Express your answer in lowest terms.

OpenStudy (anonymous):

Divide both the lengths.

OpenStudy (anonymous):

\[\large \frac{\frac{5}{6}}{\frac{3}{8}} = ??\]

OpenStudy (anonymous):

5sos?

OpenStudy (anonymous):

20/9

OpenStudy (anonymous):

@waterineyes We don't divide fractions, we multiply.. \(\dfrac{5}{6} \div \dfrac{3}{8} = \dfrac{5}{6} \times \dfrac{8}{3}\) We flip the 2nd fraction and multiply. \(\dfrac{5}{6} \times \dfrac{8}{3} \rightarrow \dfrac{5 \times 8}{6 \times 3}\) Multiply 5 * 8 and 6 * 3

OpenStudy (anonymous):

5*8=40 and 6*3=18

OpenStudy (anonymous):

You are multiplying the reverse of other fraction..

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }-\frac{ 3 }{ 8 }=\frac{ 20-9 }{ 24 }=\frac{ 11 }{ 24 }\]

OpenStudy (anonymous):

5/6 = 0.83 3/8 = 0.37 Now divide.. :P

OpenStudy (anonymous):

so 11/24 is the answer???

OpenStudy (anonymous):

And yes, we divide fractions by the process of multiplication as you have shown. :P

OpenStudy (anonymous):

who do i give the medal 2?

OpenStudy (anonymous):

To surjit..

OpenStudy (anonymous):

WAIT I HAVE ANOTHER QUESTION LAST ONE

OpenStudy (anonymous):

I still think 20/9, should be correct..

OpenStudy (anonymous):

20/9 means two times rouphly is length twice the breadth?

OpenStudy (anonymous):

Yes, so we have 40/18..

OpenStudy (anonymous):

it is simply subtract width from length.

OpenStudy (anonymous):

if the question is how many times the length to width then divide. if how much greater the length to width, then subtract.